# 3DTopia-XL: Scaling High-quality 3D Asset Generation via Primitive Diffusion

### v1

@@ Abstract
The increasing demand for high-quality 3D assets across various industries necessitates efficient and automated 3D content creation. Despite recent advancements in 3D generative models, existing methods still face challenges with optimization speed, geometric fidelity, and the lack of assets for physically based rendering (PBR). In this paper, we introduce 3DTopia-XL, a scalable native 3D generative model designed to overcome these limitations. 3DTopia-XL leverages a novel primitive-based 3D representation, PrimX, which encodes detailed shape, albedo, and material field into a compact tensorial format, facilitating the modeling of high-resolution geometry with PBR assets. On top of the novel representation, we propose a generative framework based on Diffusion Transformer (DiT), which comprises 1) Primitive Patch Compression, 2) and Latent Primitive Diffusion. 3DTopia-XL learns to generate high-quality 3D assets from textual or visual inputs. We conduct extensive qualitative and quantitative experiments to demonstrate that 3DTopia-XL significantly outperforms existing methods in generating high-quality 3D assets with fine-grained textures and materials, efficiently bridging the quality gap between generative models and real-world applications.
00footnotetext:  Corresponding author
@@ 1 Introduction
Fortunately, rapid progress has been witnessed in the field of 3D generative models recently. Existing state-of-the-art techniques can be sorted into three categories. 1) Methods based on Score Distillation Sampling (SDS)  (Poole et al., 2022; Tang et al., 2023a) lift 2D diffusion priors into 3D representation by per-scene optimization. However, these methods suffer from time-consuming optimization, poor geometry, and multifaceted inconsistency. 2) Methods based on sparse-view reconstruction (Hong et al., 2023; Xu et al., 2024a) that leverage large models to regress 3D assets from single- or multi-view images. Most of these methods are built upon triplane-NeRF (Chan et al., 2022) representation. However, due to the triplane’s parameter inefficiency, the valid parameter space is limited to low resolutions in those models, leading to relatively low-quality 3D assets. Plus, reconstruction-based models also suffer from a low-diversity problem as deterministic methods. 3) Methods as native 3D generative models (Yariv et al., 2023; Li et al., 2024c) aim to model the probabilistic distribution of 3D assets, generating 3D objects given input conditions. Yet, few of them are capable of generating high-quality 3D objects with Physically Based Rendering (PBR) assets, which are geometry, texture, and material packed into a GLB file.
@@
To address the limitations above, we propose 3DTopia-XL, a high-quality native 3D generative model for 3D assets at scale. Our key idea is scaling the powerful diffusion transformer (Peebles & Xie, 2022) on top of a novel primitive-based 3D representation. At the core of 3DTopia-XL is an efficient 3D representation, PrimX, which encodes the shape, albedo, and material of a textured mesh in a compact $N\times D$ tensor, enabling the modeling of high-resolution geometry with PBR assets. In specific, we anchor $N$ primitives to the positions sampled on the mesh surface. Each primitive is a tiny voxel, parameterized by its 3D position, a global scale factor, and corresponding spatially varied payload for SDF, RGB, and material. Note that the proposed representation differentiates itself from the shape-only representation M-SDF (Yariv et al., 2023) that PrimX encodes shape, color, and material in a unified way. It also supports efficient differentiable rendering, leading to the great potential to learn from not only 3D data but also image collections. Moreover, we carefully design initialization and fine-tuning strategy that enables PrimX to be rapidly tensorized from a textured mesh (GLB file) which is ten times faster than the triplane under the same setting.
@@
Thanks to the tensorial and compact PrimX, we scale the 3D generative modeling using latent primitive diffusion with Transformers, where we treat each 3D object as a set of primitives. In specific, the proposed 3D generation framework consists of two modules. 1) Primitive Patch Compression uses a 3D VAE for spatial compression of each individual primitive to get latent primitive tokens; and 2) Latent Primitive Diffusion leverages the Diffusion Transformers (DiT) (Peebles & Xie, 2022) to model global correlation of latent primitive tokens for generative modeling. Notably, the permutation equivariance of PrimX naturally supports training Transformers without positional encoding. The significant efficiency of the proposed representation allows us to achieve high-resolution generative training using a clean and unified framework without super-resolution or post-hoc optimization-based mesh refinement.
@@
In addition, we also carefully design algorithms for high-quality 3D PBR asset extraction from PrimX, to ensure reversible transformations between PrimX and textured mesh. An issue for most 3D generation models (Wang et al., 2024; Xu et al., 2024a) is that they use vertex coloring to represent the object’s texture, leading to a significant quality drop when exporting their generation results into mesh format. Thanks to the high-quality surface modeled by Signed Distance Field (SDF) in PrimX, we propose to extract the 3D shape with zero-level contouring and sample texture and material values in a high-resolution UV space. This leads to high-quality asset extraction with considerably fewer vertices, which is also ready to be packed into GLB format for downstream tasks.
@@ 2 Related Work
Early works on feed-forward 3D generation involves training a GAN (Goodfellow et al., 2020) from 2D image datasets (Gao et al., 2022; Chan et al., 2022; Hong et al., 2022). However, such methods fail to scale up to large-scale datasets with general 3D objects (Deitke et al., 2023b; a). Recent advancements have been focusing on feed-forward reconstruction methods that regress 3D assets from single- or multi-view images. Large Reconstruction Model (LRM) (Hong et al., 2023; He & Wang, 2023) has shown that end-to-end training of a triplane-NeRF regression model scales well to large datasets and can be highly generalizable. Although it can significantly accelerate generation speed, the generated 3D assets still exhibit relatively lower quality due to representation inefficiency and suffer from a low-diversity problem as a deterministic method. Subsequent works have extended this method to improve generation quality. For example, using multi-view images (Xu et al., 2024a; Li et al., 2023; Wang et al., 2023a; Siddiqui et al., 2024; Xie et al., 2024; Wang et al., 2024; Jiang et al., 2024; Boss et al., 2024) generated by 2D diffusion models as the input can effectively enhance the visual quality. However, the generative capability is actually enabled by the frontend multi-view diffusion models (Shi et al., 2023; Li et al., 2024b; Long et al., 2023) which cannot produce multi-view images with accurate 3D consistency. Another direction is to use more efficient 3D representations such as Gaussian Splatting (Kerbl et al., 2023; Tang et al., 2024; Xu et al., 2024c; Zhang et al., 2024b; Yi et al., 2024; Chen et al., 2024) and triangular mesh (Zhang et al., 2024a; Li et al., 2024a; Wei et al., 2024; Zou et al., 2023). However, few of them can generate high-quality PBR assets with sampling diversity.
3D Diffusion Generative Models.
@@
Similar to 2D diffusion models for image generation, efforts have been made to train a 3D native diffusion model on conditional 3D generation. However, unlike the universal image representation in 2D, there are many different choices for 3D representations. Voxel-based methods (Müller et al., 2023) can be directly extended from 2D methods, but they are constrained by the demanding memory usage, and suffer from scaling up to high-resolution data. Point cloud based methods (Nichol et al., 2022; Nash & Williams, 2017) are memory-efficient and can adapt to large-scale datasets, but they hardly represent the watertight and solid surface of the 3D assets. Implicit representations such as triplane-NeRF offer a better balance between memory and quality (Jun & Nichol, 2023; Gupta et al., 2023; Cheng et al., 2023; Ntavelis et al., 2023; Cao et al., 2023; Chen et al., 2023a; Wang et al., 2023b; Liu et al., 2023a) There are also methods based on other representations such as meshes and primitives (Liu et al., 2023b; Yariv et al., 2023; Chen et al., 2023b; Xu et al., 2024b; Yan et al., 2024). However, these methods still struggle with generalization or producing high-quality assets. Recent methods attempt to adapt latent diffusion models to 3D (Zhang et al., 2023; Zhao et al., 2023; Zhang et al., 2024c; Wu et al., 2024; Li et al., 2024c; Lan et al., 2024; Hong et al., 2024; Tang et al., 2023b). These methods first train a 3D compression model such as a VAE to encode 3D assets into a more compact form, which allows the diffusion model to train more effectively and show strong generalization. However, they either suffer from low-resolution results or are incapable of modeling PBR materials. In this paper, we propose a new 3D latent diffusion model based on a novel representation, PrimX, which can be efficiently computed from a textured mesh and unpacked into high-resolution geometry with PBR materials.
@@ 3.1 PrimX: An Efficient Representation for Shape, Texture, and Material
Before diving into details, we outline the following design principles for 3D representation in the context of high-quality large-scale 3D generative models: 1) Parameter-efficient: provides a good trade-off between approximation error and parameter count; 2) Rapidly tensorizable: can be efficiently transformed into a tensorial structure, which facilitates generative modeling with modern neural architectures; 3) Differentiably renderable: compatible with differentiable renderer, enabling learning from both 3D and 2D data.
Given the aforementioned principles, we propose a novel primitive-based 3D representation, namely PrimX, which represents the 3D shape, texture, and material of a textured mesh as a compact $N\times D$ tensor. It can be efficiently computed from a textured mesh (typically a GLB file) and directly rendered into 2D images via a differentiable rasterizer.
@@ 3.1.1 Definition
Preliminaries.
Given a textured 3D mesh, we denote its 3D shape as ${\mathcal{S}}\in{\mathbb{R}}^{3}$, where ${\mathbf{x}}\in{\mathcal{S}}$ are spatial points inside the occupancy of the shape, and ${\mathbf{x}}\in\partial{\mathcal{S}}$ are the points on the shape’s boundary, i.e., the shape’s surface. We model the 3D shape as SDF as follows:
$F_{{\mathcal{S}}}^{\mathrm{SDF}}({\mathbf{x}})=\begin{cases}-d({\mathbf{x}},\partial{\mathcal{S}}),\;\;{\mathbf{x}}\in{\mathcal{S}}\\ d({\mathbf{x}},\partial{\mathcal{S}}),\;\;\mathrm{elsewise}\end{cases}\;\;\mathrm{s.t.}\;\;d({\mathbf{x}},\partial{\mathcal{S}})=\min_{{\mathbf{y}}\in{\mathcal{S}}}||{\mathbf{x}}-{\mathbf{y}}||_{2}.$ (1)
Moreover, given the neighborhood of shape surface, ${\mathcal{U}}(\partial{\mathcal{S}},\delta)=\{d({\mathbf{x}},\partial{\mathcal{S}})<\delta\}$, the space-varied color function and material function of the target mesh are defined by:
$F_{{\mathcal{S}}}^{\mathrm{RGB}}({\mathbf{x}})=\begin{cases}C({\mathbf{x}}),\;\;{\mathbf{x}}\in{\mathcal{U}}\\ 0,\;\;\mathrm{elsewise}\end{cases}F_{{\mathcal{S}}}^{\mathrm{Mat}}({\mathbf{x}})=\begin{cases}\rho({\mathbf{x}}),\;\;{\mathbf{x}}\in{\mathcal{U}}\\ 0,\;\;\mathrm{elsewise}\end{cases},$ (2)
where $C({\mathbf{x}}):{\mathbb{R}}^{3}\xrightarrow{}{\mathbb{R}}^{3}$ and $\rho({\mathbf{x}}):{\mathbb{R}}^{3}\xrightarrow{}{\mathbb{R}}^{2}$ are corresponding texture sampling functions to get albedo and material (metallic and roughness) from UV-aligned texture maps given the 3D point ${\mathbf{x}}$. Eventually, all shape, texture, and material information of a 3D mesh can be parameterized by the volumetric function $F_{{\mathcal{S}}}=(F_{{\mathcal{S}}}^{\mathrm{SDF}}\oplus F_{{\mathcal{S}}}^{\mathrm{RGB}}\oplus F_{{\mathcal{S}}}^{\mathrm{Mat}}):{\mathbb{R}}^{3}\xrightarrow{}{\mathbb{R}}^{6}$, where $\oplus$ denotes concatenation.
PrimX Representation.
We aim to approximate $F_{{\mathcal{S}}}$ with a neural volumetric function $F_{{\mathcal{V}}}:{\mathbb{R}}^{3}\xrightarrow{}{\mathbb{R}}^{6}$ parameterized by a $N\times D$ tensor ${\mathcal{V}}$. For efficiency, our key insight is to define $F_{{\mathcal{V}}}$ as a set of $N$ volumetric primitives distributed on the surface of the mesh:
@@
Inspired by Yariv et al. (2023) where mosaic voxels are globally weighted to get a smooth surface, the approximation of a textured mesh is then defined as a weighted combination of primitives:
$F_{{\mathcal{V}}}({\mathbf{x}})=\sum_{k=1}^{N}w_{k}({\mathbf{x}}){\mathcal{I}}({\mathcal{V}}_{k},\frac{{\mathbf{x}}-{\mathbf{t}}_{k}}{s_{k}}),$ (4)
@@
$w_{k}({\mathbf{x}})=\frac{\hat{w}_{k}({\mathbf{x}})}{\sum_{j=1}^{N}\hat{w}_{j}({\mathbf{x}})},\;\;\mathrm{s.t.}\;\;\hat{w}_{k}({\mathbf{x}})=\max(0,1-||\frac{{\mathbf{x}}-{\mathbf{t}}_{k}}{s_{k}}||_{\infty}).$ (5)
@@
where we use an exponent of the SDF field to represent the opacity field. And $\alpha$ is the hyperparameter that controls the variance of the opacity field during this conversion. where we set $\alpha=0.005$.
To wrap up, the learnable parameters of a textured 3D mesh modeled by PrimX are primitive position ${\mathbf{t}}\in{\mathbb{R}}^{N\times 3}$, primitive scale $s\in{\mathbb{R}}^{N\times 1}$, and voxel payload ${\bm{X}}\in{\mathbb{R}}^{N\times a^{3}\times 6}$ for SDF, color, and material. Therefore, each textured mesh can be represented as a compact $N\times D$ tensor, where $D=3+1+a^{3}\times 6$ by concatenation.
PBR Asset Extraction.
Once PrimX is constructed, it encodes all geometry and appearance information of the target mesh within the $N\times D$. Now, we introduce our efficient algorithm to convert PrimX back into a textured mesh in GLB file format. For geometry, we can easily extract the corresponding 3D shape with Marching Cubes algorithm on zero level set of $F_{{\mathcal{V}}}^{\mathrm{SDF}}$. For PBR texture maps, we first perform UV unwrapping using xatlas in a 1k resolution UV space. Then, we get sampling points in 3D using nvdiffrast and query $F_{{\mathcal{V}}}^{\mathrm{RGB}},F_{{\mathcal{V}}}^{\mathrm{Mat}}$ to get corresponding albedo and material values. Note that, we mask the UV space to get the index of valid vertices for efficient queries. Moreover, we dilate the UV texture maps and inpaint the dilated region with the nearest neighbors of existing textures, ensuring albedo and material maps smoothly blend outwards for anti-aliasing. Finally, we pack geometry, UV mapping, albedo, metallic, and roughness maps into a GLB file, which is ready for the graphics engine and various downstream tasks.

### v2

@@ Abstract
The increasing demand for high-quality 3D assets across various industries necessitates efficient and automated 3D content creation. Despite recent advancements in 3D generative models, existing methods still face challenges with optimization speed, geometric fidelity, and the lack of assets for physically based rendering (PBR). In this paper, we introduce 3DTopia-XL, a scalable native 3D generative model designed to overcome these limitations. 3DTopia-XL leverages a novel primitive-based 3D representation, PrimX, which encodes detailed shape, albedo, and material field into a compact tensorial format, facilitating the modeling of high-resolution geometry with PBR assets. On top of the novel representation, we propose a generative framework based on Diffusion Transformer (DiT), which comprises 1) Primitive Patch Compression, 2) and Latent Primitive Diffusion. 3DTopia-XL learns to generate high-quality 3D assets from textual or visual inputs. Extensive qualitative and quantitative evaluations are conducted to demonstrate that 3DTopia-XL significantly outperforms existing methods in generating high-quality 3D assets with fine-grained textures and materials, efficiently bridging the quality gap between generative models and real-world applications.
@@ 1 Introduction
Fortunately, rapid progress has been witnessed in the field of 3D generative models recently. Existing state-of-the-art techniques can be sorted into three categories. 1) Methods based on Score Distillation Sampling (SDS)  [46, 55] lift 2D diffusion priors into 3D representation by per-scene optimization. However, these methods suffer from time-consuming optimization, poor geometry, and multifaceted inconsistency. 2) Methods based on sparse-view reconstruction [19, 68] that leverage large models to regress 3D assets from single- or multi-view images. Most of these methods are built upon triplane-NeRF [5] representation. However, due to the triplane’s parameter inefficiency, the valid parameter space is limited to low resolutions in those models, leading to relatively low-quality 3D assets. Plus, reconstruction-based models also suffer from a low-diversity problem as deterministic methods. 3) Methods as native 3D generative models [72, 30] aim to model the probabilistic distribution of 3D assets, generating 3D objects given input conditions. Yet, few of them are capable of generating high-quality 3D objects with Physically Based Rendering (PBR) assets, which are geometry, texture, and material packed into a GLB file.
@@
To address the limitations above, we propose 3DTopia-XL, a high-quality native 3D generative model for 3D assets at scale. Our key idea is scaling the powerful diffusion transformer [45] on top of a novel primitive-based 3D representation. At the core of 3DTopia-XL is an efficient 3D representation, PrimX, which encodes the shape, albedo, and material of a textured mesh in a compact $N\times D$ tensor, enabling the modeling of high-resolution geometry with PBR assets. In specific, we anchor $N$ primitives to the positions sampled on the mesh surface. Each primitive is a tiny voxel, parameterized by its 3D position, a global scale factor, and corresponding spatially varied payload for SDF, RGB, and material. Note that the proposed representation differentiates itself from the shape-only representation M-SDF [72] that PrimX encodes shape, color, and material in a unified way. It also supports efficient differentiable rendering, leading to the great potential to learn from not only 3D data but also image collections. Moreover, we carefully design initialization and fine-tuning strategy that enables PrimX to be rapidly tensorized from a textured mesh (GLB file) which is ten times faster than the triplane under the same setting.
@@
Thanks to the tensorial and compact PrimX, we scale the 3D generative modeling using latent primitive diffusion with Transformers, where we treat each 3D object as a set of primitives. In specific, the proposed 3D generation framework consists of two modules. 1) Primitive Patch Compression uses a 3D VAE for spatial compression of each individual primitive to get latent primitive tokens; and 2) Latent Primitive Diffusion leverages the Diffusion Transformers (DiT) [45] to model global correlation of latent primitive tokens for generative modeling. The significant efficiency of the proposed representation allows us to achieve high-resolution generative training using a clean and unified framework without super-resolution to upscale the underlying 3D representation or post-hoc optimization-based mesh refinement.
@@
In addition, we also carefully design algorithms for high-quality 3D PBR asset extraction from PrimX, to ensure reversible transformations between PrimX and textured mesh. An issue for most 3D generation models [62, 68] is that they use vertex coloring to represent the object’s texture, leading to a significant quality drop when exporting their generation results into mesh format. Thanks to the high-quality surface modeled by Signed Distance Field (SDF) in PrimX, we propose to extract the 3D shape with zero-level contouring and sample texture and material values in a high-resolution UV space. This leads to high-quality asset extraction with considerably fewer vertices, which is also ready to be packed into GLB format for downstream tasks.
@@ 2 Related Work
Sparse-view Reconstruction Models. Recent advancements have been focusing on deterministic reconstruction methods that regress 3D assets from single- or multi-view images. Large Reconstruction Model (LRM) [19, 14] has shown that end-to-end training of a triplane-NeRF [5] regression model scales well to large datasets and can be highly generalizable. Although it can significantly accelerate generation speed, the generated 3D assets still exhibit relatively lower quality due to representation inefficiency and suffer from a low-diversity problem as a deterministic method. Subsequent works have extended this method to improve generation quality. For example, using multi-view images [68, 27, 60, 53, 66, 62, 20, 2] generated by 2D diffusion models as the input can effectively enhance the visual quality. However, the generative capability is actually enabled by the frontend multi-view diffusion models [52, 29, 36, 32] which cannot produce multi-view images with accurate 3D consistency. Another direction is to use more efficient 3D representations such as Gaussian Splatting [23, 56, 70, 76, 73, 6] and triangular mesh [75, 28, 63, 80, 34]. However, few of them can generate high-quality PBR assets with sampling diversity.
@@
Native 3D Diffusion Models. Similar to 2D diffusion models for image generation, efforts have been made to train a 3D native diffusion model [31, 79, 49, 9] on conditional 3D generation. However, unlike the universal image representation in 2D, there are many different choices for 3D representations. Voxel-based methods [40] can be directly extended from 2D methods, but they are constrained by the demanding memory usage, and suffer from scaling up to high-resolution data. Point cloud based methods [42, 41] are memory-efficient and can adapt to large-scale datasets, but they hardly represent the watertight and solid surface of the 3D assets. Implicit representations such as triplane-NeRF offer a better balance between memory and quality [21, 13, 9, 43, 4, 7, 61, 33]. There are also methods based on other representations such as meshes and primitives [35, 72, 8, 69, 71]. However, these methods still struggle with generalization or producing high-quality assets. Recent methods attempt to adapt latent diffusion models to 3D [74, 78, 77, 65, 30, 26, 18, 57]. These methods first train a 3D compression model such as a VAE to encode 3D assets into a more compact form, which allows the diffusion model to train more effectively and show strong generalization. However, they either suffer from low-resolution results or are incapable of modeling PBR materials. In this paper, we propose a new 3D latent diffusion model on a novel representation, PrimX, which can be efficiently computed from a textured mesh and unpacked into high-resolution geometry with PBR materials.
@@ 3.1 PrimX: An Efficient Representation for Shape, Texture, and Material
Before diving into details, we outline the following design principles for 3D representation in the context of high-quality large-scale 3D generative models: 1) Parameter-efficient: provides a good trade-off between approximation error and parameter count; 2) Rapidly tensorizable: can be efficiently transformed into a tensor, which facilitates generative modeling with modern neural architectures; 3) Differentiably renderable: compatible with differentiable renderer, enabling learning from both 3D and 2D data.
Given the aforementioned principles, we propose a novel primitive-based 3D representation, namely PrimX, which represents the 3D shape, texture, and material of a textured mesh as a compact $N\times D$ tensor. It can be efficiently computed from a textured mesh (in GLB format) and directly rendered into 2D images via a differentiable rasterizer.
@@ 3.1.1 Definition
Preliminaries. Given a textured 3D mesh, we denote its 3D shape as ${\mathcal{S}}\in{\mathbb{R}}^{3}$, where ${\mathbf{x}}\in{\mathcal{S}}$ are spatial points inside the occupancy of the shape, and ${\mathbf{x}}\in\partial{\mathcal{S}}$ are the points on the shape’s boundary, i.e. the shape’s surface. We model the 3D shape as SDF as follows:
$F_{{\mathcal{S}}}^{\mathrm{SDF}}({\mathbf{x}})=\operatorname{sign}({\mathbf{x}},{\mathcal{S}})\cdot d({\mathbf{x}},\partial{\mathcal{S}}),$ (1)
where $d({\mathbf{x}},\partial{\mathcal{S}})=\min_{{\mathbf{y}}\in{\mathcal{S}}}||{\mathbf{x}}-{\mathbf{y}}||_{2}$ is the distance function, $\operatorname{sign}({\mathbf{x}},{\mathcal{S}})=1$ when ${\mathbf{x}}\in{\mathcal{S}}$, and $\operatorname{sign}({\mathbf{x}},{\mathcal{S}})=-1$ when ${\mathbf{x}}\notin{\mathcal{S}}$. Moreover, given the neighborhood of the surface, ${\mathcal{U}}(\partial{\mathcal{S}},\delta)=\{d({\mathbf{x}},\partial{\mathcal{S}})<\delta\}$, the space-varied color and material functions of the target mesh are defined:
$F_{{\mathcal{S}}}^{\mathrm{RGB}}({\mathbf{x}})=\operatorname{sign}({\mathbf{x}},{\mathcal{U}})\cdot C({\mathbf{x}}),F_{{\mathcal{S}}}^{\mathrm{Mat}}({\mathbf{x}})=\operatorname{sign}({\mathbf{x}},{\mathcal{U}})\cdot\rho({\mathbf{x}}),$ (2)
where $\operatorname{sign}({\mathbf{x}},{\mathcal{U}})=1$ when ${\mathbf{x}}\in{\mathcal{U}}$, and $\operatorname{sign}({\mathbf{x}},{\mathcal{U}})=0$ when ${\mathbf{x}}\notin{\mathcal{U}}$. The $C({\mathbf{x}}):{\mathbb{R}}^{3}\xrightarrow{}{\mathbb{R}}^{3}$ and $\rho({\mathbf{x}}):{\mathbb{R}}^{3}\xrightarrow{}{\mathbb{R}}^{2}$ are corresponding texture sampling functions to get albedo and material (metallic and roughness) from UV-aligned texture maps given the 3D point ${\mathbf{x}}$. Eventually, all shape, texture, and material information of a 3D mesh can be parameterized by the volumetric function $F_{{\mathcal{S}}}=(F_{{\mathcal{S}}}^{\mathrm{SDF}}\oplus F_{{\mathcal{S}}}^{\mathrm{RGB}}\oplus F_{{\mathcal{S}}}^{\mathrm{Mat}}):{\mathbb{R}}^{3}\xrightarrow{}{\mathbb{R}}^{6}$, where $\oplus$ denotes concatenation.
PrimX Representation. We aim to approximate $F_{{\mathcal{S}}}$ with a neural volumetric function $F_{{\mathcal{V}}}:{\mathbb{R}}^{3}\xrightarrow{}{\mathbb{R}}^{6}$ parameterized by a $N\times D$ tensor ${\mathcal{V}}$. For efficiency, our key insight is to define $F_{{\mathcal{V}}}$ as a set of $N$ volumetric primitives distributed on the surface of the mesh:
@@
Inspired by Yariv et al. [72] where mosaic voxels are globally weighted to get a smooth surface, the approximation of a textured mesh is then defined as a weighted combination of primitives:
$F_{{\mathcal{V}}}({\mathbf{x}})=\sum\nolimits_{k=1}^{N}[w_{k}({\mathbf{x}})\cdot{\mathcal{I}}({\bm{X}}_{k},({\mathbf{x}}-{\mathbf{t}}_{k})/{s_{k}})],$ (4)
@@
$w_{k}({\mathbf{x}})=\frac{\hat{w}_{k}({\mathbf{x}})}{\sum_{j=1}^{N}\hat{w}_{j}({\mathbf{x}})},$ (5)
$\mathrm{s.t.}\;\;\hat{w}_{k}({\mathbf{x}})=\max(0,1-||\frac{{\mathbf{x}}-{\mathbf{t}}_{k}}{s_{k}}||_{\infty}).$ (6)
@@
where $T(t)$ is an exponential function of the SDF field to represent the opacity field:
[…]
And $\alpha=0.005$ is the hyperparameter that controls the variance of the opacity field during this conversion.
To wrap up, the learnable parameters of a textured 3D mesh modeled by PrimX are primitive position ${\mathbf{t}}\in{\mathbb{R}}^{N\times 3}$, primitive scale $s\in{\mathbb{R}}^{N\times 1}$, and voxel payload ${\bm{X}}\in{\mathbb{R}}^{N\times a^{3}\times 6}$ for SDF, albedo, and material. Therefore, each textured mesh can be represented as a compact $N\times D$ tensor, where $D=3+1+a^{3}\times 6$ by concatenation.
PBR Asset Extraction. Once PrimX is constructed, it encodes all geometry and appearance information of the target mesh within the $N\times D$ tensor. Now, we introduce our efficient algorithm to convert PrimX back into a textured mesh in GLB file format. For geometry, we can easily extract the corresponding 3D shape with Marching Cubes algorithm [37] on zero level set of $F_{{\mathcal{V}}}^{\mathrm{SDF}}$. For PBR texture maps, we first perform UV unwrapping in a high-resolution UV space ($1024\times 1024$). Then, we get sampling points in 3D and query $\{F_{{\mathcal{V}}}^{\mathrm{RGB}},F_{{\mathcal{V}}}^{\mathrm{Mat}}\}$ to get corresponding albedo and material values. Note that, we mask the UV space to get the index of valid vertices for efficient queries. Moreover, we dilate the UV texture maps and inpaint the dilated region with the nearest neighbors of existing textures, ensuring albedo and material maps smoothly blend outwards for anti-aliasing. Finally, we pack geometry, UV mapping, albedo, and material maps into a GLB file, which is ready for the graphics engine and various downstream tasks.